OpenStudy (agl202):

;) What is the simplified form of 1 over x minus 2 over x squared plus x ? x minus 1 over the quantity x times x plus 1 1 minus x over the quantity x times x plus 1 3 minus x over the quantity x times x minus 1 x plus 2 over the quantity x times x minus 1 I am positive it's A.

OpenStudy (md152727):

whatgrade are you?

OpenStudy (agl202):

@AlexandervonHumboldt2 Pretty plz?

OpenStudy (alexandervonhumboldt2):

sureeeeeeeeeeeeeeeeeeeeeeeeeeeeee

OpenStudy (md152727):

what grade is this?

OpenStudy (alexandervonhumboldt2):

\[\frac{ 1 }{ x }-\frac{ 2 }{ x^2+x }?\]

OpenStudy (alexandervonhumboldt2):

ok so \[\frac{ x+1 }{ x^2+x }-\frac{ 2 }{ x^2+x }==\frac{ x-1 }{ x^2+x }=\frac{ x-1 }{ x*(x+1) }\]

OpenStudy (alexandervonhumboldt2):

so yeah A seems to be correct

OpenStudy (alexandervonhumboldt2):

not sure that this is a Grad question lol

OpenStudy (agl202):

Thx. It's alg2. lol

OpenStudy (agl202):

alg=agl

OpenStudy (alexandervonhumboldt2):

lol

OpenStudy (alexandervonhumboldt2):

е е е тра ли ва ли

OpenStudy (agl202):

^ O.o

OpenStudy (alexandervonhumboldt2):

O_o